A newly discovered metalloprotein, which contains
Answer #1

atomic mass of Mn = 55

total mass of Mn in molecule of protein = 3 * 55 = 165

1 mole of protein contains 165 g of Mn then

165 g is 0.818% of protein then molar mass of protein = 165*100/0.818

molar mass of protein = 20171.15 g/mol